OpenStudy (anonymous):

What is the exact value of 5 times z all over the square root of z to the fifth power.? A:5 the square root of z all over z cubed. B:5 the square root of z all over z squared. C:5 the square root of z all over z to the fourth power. D:5 the square root of z all over z.

OpenStudy (campbell_st):

\[\frac{5z}{\sqrt{z^5}} = 5z \times z^{-5/2}\] when multiplying add the powers